# Parameter control for eccentric, precessing binary black hole simulations with SpEC

### v1

@@ IV.2 Effect of mass ratio
Figure 9: Effect of mass ratio on parameter convergence. We plot the difference between the target and the fitted eccentricity (left), semi-major axis (middle), and spin angle (right) for different values of the mass ratio (linestyles) and direction of the target in-plane spin, $\arccos(\vec{\chi}_{x}\cdot\vec{\chi}_{T})$ (colorbar). We choose mass ratios $q=1$, (solid), $q=2$, (dashed), and $q=3$, (dotted) and constant target spin magnitude $|\vec{\chi}_{1,T}|=0.3$. In all cases, the parameters converge after $\leq 5$ iterations.
@@ V Discussion
We have devised a method to achieve numerical BBH simulations in SpEC with target orbital parameters and BH spins beginning at a reference time, $t=t_{T}$. Our method is an extension of previous efforts to target vanishing eccentricity. It is based on iteratively performing short numerical simulations, fitting the orbit, and calculating updated initial data parameters until a tolerance threshold is achieved. We have confirmed convergence for the orbital eccentricity and the spin directions to tolerances of $\mathcal{O}(10^{-3})$ and $\mathcal{O}({\rm deg})$ respectively. We have also shown that this approach is effective for binaries with mass ratios $q\leq 3$, eccentricities $e\leq 0.65$, and spin magnitudes $\chi\leq 0.75$ in $\leq 7$ iterations. Overall, convergence is achieved with fewer iterations for smaller spin magnitudes and eccentricities, while the mass ratio has minimal impact on the rate of convergence.
The number of iterations required sensitively depends on the initial parameters of the first iteration. In the absence of a better solution, we calculate them based on the target orbital parameters and the 1PN equations of motion. A more robust procedure to calculate the initial data parameters of the first iteration would improve convergence. For example, rough initial data parameters for quasicircular simulations can be obtained for SpEC based on a Gaussian process regression fitted on the existing extensive catalog of quasicircular simulations Boyle et al. (2019). Once an appropriately large eccentric catalog is available, similar techniques could be adopted for calculating input parameters of future eccentric simulations. Other approaches include using higher-PN order or effective-one-body equations of motion Nagar et al. (2021).
Currently and inspired by the quasicircular case, the tolerance threshold for terminating the parameter control loop is an accuracy of $7\times 10^{-4}$ in eccentricity while the other parameters are ignored. In practice, this choice still achieves good convergence for all parameters, notably the semi-major axis and spin directions. In general, the appropriate tolerance threshold depends on the application. If future applications, such as building surrogate models that require simulations in sparsely covered regions of the parameter space Varma et al. (2019b), require greater precision in other parameters, adjusting the tolerance threshold in the parameter control loop is trivial.
@@ Acknowledgements.
The computations presented here were conducted partly in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. Computations for this work were also partly performed with the Wheeler cluster at Caltech. We thank Saul Teukolsky for discussions on fitting methods and the SXS group for advice on using SpEC. This work was supported by the Sherman Fairchild Foundation, and by NSF Grants No. PHY-2309211, PHY-2309231, and OAC-2209656 at Caltech.
@@ A.1 Eccentric post-Newtonian dynamics
We characterize a generic Keplerian orbit at Newtonian order (0PN) by its semi-major axis $a$, eccentricity $e$, and mean anomaly $\ell$. We relate these three quantities through Kepler’s equations: $\Omega t=u(t)-e\sin{u(t)}$ and $\Omega=\sqrt{M/a^{3}}$. We can invert Kepler’s equations to obtain the eccentric anomaly, $u(t)$:
$u(t)\equiv 2\tan^{-1}{[\sqrt{(1-e)/(1+e)}\tan{(l/2})]}\,.$ (15)
An approximation of this expression to linear order is $u(t)=\Omega t+\ell$. Furthermore, we define the separation of the two bodies in orbit as: $r(t)=a(1-e\cos u(t))$.
Next, we introduce 1PN corrections to the semi-major axis, $a$, and eccentricity, $e$. We denote 1PN-corrected quantities with a bar ($\bar{a}$, $\bar{e}$, etc.). We relate the 1PN-corrected quantities to Newtonian quantities via:
@@
$u(t)=\frac{1}{a^{3/2}}\left(1-\frac{9-\eta}{2}\frac{1}{a}\right)t+\ell\,,$ (19)
where the 1PN correction depends on the symmetric mass ratio $\eta$.
In post-Newtonian theory, noncircularity is described with multiple eccentricity parameters: the radial $\tilde{e}$, temporal $e_{t}$, and angular eccentricity $e_{\phi}$. We can similarly define another set of 1PN quantities in terms of the radial eccentricity parameter, $\tilde{e}$. This second set of 1PN quantities are denoted by a tilde ($\tilde{a}$, $\tilde{e}$, etc.). We can relate this set of 1PN-corrected quantities to the Keplerian orbital parameters via:
@@
Again, we preserve the separation relationship: $\tilde{r}(t)=\tilde{a}(1-\tilde{e}\cos u(t))$. The remaining two PN eccentricity parameters are defined in terms of the Newtonian-order eccentricity:
@@
The derivation of $\dot{\Omega}_{e}$ follows Poisson and Will Poisson and Will . We recall that $\eta$ is the symmetric mass ratio, $a$ is the Keplerian semi-major axis, and $e$ is the Keplerian eccentricity. We then introduce the 1PN-corrected orbital quantities to supplement those from Sec. A.1:
$P$ $=2\pi\sqrt{\frac{a^{3}}{M}}\left[1+\frac{1}{2}(9-\eta)\frac{M}{a}\right]\,,$ (25)
@@
$=\frac{2\pi}{P(1-e_{t}\cos u(t))}\frac{\tilde{h}}{\tilde{a}^{2}}\frac{-2\tilde{e}\sin u(t)(1-\tilde{e}\cos u(t))}{(1-\tilde{e}\cos u(t))^{4}}\,,$ (27)
where we multiplied by $1=(1-\tilde{e}\cos u(t))/(1-\tilde{e}\cos u(t))$ to get an even power of $(1-\tilde{e}\cos u(t))$ in the denominator. To this 1PN order, we apply $(1-\tilde{e}\cos u(t))^{2}=(1-e_{t}\cos u(t))(1-e_{\phi}\cos u(t))$. Thus, Eq. (27) simplifies to
@@ A.3 Deriving the mapping from $\theta_{\rm orb}$ to $\theta_{\rm ID}$
Again, we follow Ref. Poisson and Will for deriving a mapping between the orbital parameters $\vec{\theta}_{\rm orb}=(a,e,\ell)$ and the initial data parameters $\vec{\theta}_{\rm ID}=(\Omega,v_{r},D)$.
@@
The $a$ and $e$ in Eq. (31) are the Newtonian parameters for semi-major axis and eccentricity, respectively.
@@
Finally, we consider the angular velocity, $\Omega(t)$. We begin with the first line of Eq. (27) for the angular velocity:
@@
Then, we substitute Eqs. (20) and (26) to achieve:
@@
Further substituting the 1PN approximations of Eqs. (26), (20), (21) yields the final expression:
@@ A.4 Spin update derivations
First, we define $R_{\rm BH\to in}(t=t_{T})$, the rotation from the co-rotating BH frame into the inertial frame. The target spins $\vec{\chi}_{A,T}(t=t_{T})$ are defined in the BH frame which might not coincide with the inertial frame at $t_{T}$. The rotation matrix $R_{\rm BH\to in}(t=t_{T})$ which transforms from the BH to the inertial frame at $t_{T}$, is formed through the basis vectors $\hat{n}(t_{T}),\hat{\lambda}(t_{T})$, and $\hat{L}(t_{T})$, defined in the inertial frame. Again, we utilize the inertial frame vector notation from Ref. Buonanno et al. (2011):

### v2

@@ V Discussion
We have devised a method to achieve numerical BBH simulations in SpEC with target orbital parameters and BH spins beginning at a reference time, $t=t_{T}$. Our method is an extension of previous efforts to target vanishing eccentricity. It is based on iteratively performing short numerical simulations, fitting the orbit, and calculating updated initial data parameters until a tolerance threshold is achieved. We have confirmed convergence for the orbital eccentricity and the spin directions to tolerances of $\mathcal{O}(10^{-3})$ and $\mathcal{O}({\rm deg})$ respectively. We have also shown that this approach is effective for binaries with mass ratios $q\leq 3$, eccentricities $e\leq 0.65$, and spin magnitudes $\chi\leq 0.75$ in $\leq 7$ iterations.
Overall, convergence is achieved with fewer iterations for smaller spin magnitudes and eccentricities, while the mass ratio has minimal impact on the rate of convergence. [^2] Each iteration’s computational time and resources depend on the target orbital parameters. For moderate spin magnitudes and comparable masses at moderate eccentricity, one iteration takes only a few hours.  For large eccentricity (where the orbital period is longer), or for large spin magnitudes and/or mass-ratio (where the code runs more slowly), the time per iteration can reach several days.  Furthermore, the first few iterations of the parameter control are run at a coarser resolution. When the eccentricity is close to the target, SpEC switches to a higher resolution. This reduces computational cost for the first few iterations of the parameter control. The number of iterations required sensitively depends on the initial parameters of the first iteration. In the absence of a better solution, we calculate them based on the target orbital parameters and the 1PN equations of motion. A more robust procedure to calculate the initial data parameters of the first iteration would improve convergence. For example, rough initial data parameters for quasicircular simulations can be obtained for SpEC based on a Gaussian process regression fitted on the existing extensive catalog of quasicircular simulations Boyle et al. (2019). Once an appropriately large eccentric catalog is available, similar techniques could be adopted for calculating input parameters of future eccentric simulations. Other approaches include using higher-PN order or effective-one-body equations of motion Nagar et al. (2021).
Currently and inspired by the quasicircular case, the tolerance threshold for terminating the parameter control loop is an accuracy of $7\times 10^{-4}$ in eccentricity while the other parameters are ignored. In practice, this choice still achieves good convergence for all parameters, notably the semimajor axis and spin directions. In general, the appropriate tolerance threshold depends on the application. If future applications, such as building surrogate models that require simulations in sparsely covered regions of the parameter space Varma et al. (2019b), require greater precision in other parameters, adjusting the tolerance threshold in the parameter control loop is trivial.
@@
Figure 9: Effect of mass ratio on parameter convergence. We plot the difference between the target and the fitted eccentricity (left), semimajor axis (middle), and spin angle (right) for different values of the mass ratio (line styles) and direction of the target in-plane spin, $\arccos(\vec{\chi}_{x}\cdot\vec{\chi}_{T})$ (color bar). We choose mass ratios $q=1$, (solid), $q=2$, (dashed), and $q=3$, (dotted) and constant target spin magnitude $|\vec{\chi}_{1,T}|=0.3$. In all cases, the parameters converge after $\leq 5$ iterations.
@@ Acknowledgements.
The computations presented here were conducted partly in the Resnick High Performance Computing Center, a facility supported by Resnick Sustainability Institute at the California Institute of Technology. Computations for this work were also partly performed with the Wheeler cluster at Caltech. We thank Saul Teukolsky for discussions on fitting methods and the SXS group for advice on using SpEC. This work was supported by the Sherman Fairchild Foundation, and by NSF Grants No. PHY-2309211, No. PHY-2309231, and No. OAC-2209656 at Caltech.
Data Availability
The data that support the findings of this article are not publicly available because they are owned by a third party and the terms of use prevent public distribution. The data are available from the authors upon reasonable request.
@@ A.1 Eccentric post-Newtonian dynamics
We characterize a generic Keplerian orbit at Newtonian order (0PN) by its semimajor axis $a$, eccentricity $e$, and mean anomaly $\ell$. We relate these three quantities through Kepler’s equations
$\bar{\Omega}t=u(t)-e\sin{u(t)}-\ell$ (15)
and $\bar{\Omega}=\sqrt{M/a^{3}}$. We define $\ell$ as the mean anomaly at the epoch $t=0$.
An approximation of this expression for small eccentricity is $u(t)=\bar{\Omega}t+\ell$. Furthermore, we define the separation of the two bodies in orbit as: $r(t)=a(1-e\cos u(t))$.
Next, we introduce 1PN corrections to the semimajor axis, $a$, and eccentricity, $e$. We denote 1PN-corrected quantities with a bar ($\bar{a}$, $\bar{e}$, etc.). We relate the 1PN-corrected quantities to Newtonian quantities via
@@
$u(t)=\sqrt{\frac{M}{a^{3}}}\left[1-\frac{(9-\eta)}{2}\frac{M}{a}\right]t+\ell\,,$ (19)
where the 1PN correction depends on the symmetric mass ratio $\eta$. Here, we have assumed a small eccentricity in Eq. (15) such that $e\sin u(t)$ goes to zero. Then we insert the 1PN expression for $\bar{\Omega}$. Although numerically inverting Eq. (15) is more accurate for larger eccentricities, it is more computationally expensive than utilizing Eq. (19).
In post-Newtonian theory, noncircularity is described with multiple eccentricity parameters: the radial $\tilde{e}$, temporal $e_{t}$, and angular eccentricity $e_{\phi}$. We can similarly define another set of 1PN quantities in terms of the radial eccentricity parameter, $\tilde{e}$. This second set of 1PN quantities are denoted by a tilde ($\tilde{a}$, $\tilde{e}$, etc.). We can relate this set of 1PN-corrected quantities to the Keplerian orbital parameters via
@@
Again, we preserve the separation relationship: $\tilde{r}(t)=\tilde{a}(1-\tilde{e}\cos u(t))$. The remaining two PN eccentricity parameters are defined in terms of the Newtonian-order eccentricity
@@
The derivation of $\dot{\Omega}_{e}$ follows Poisson and Will Poisson and Will (2014). We recall that $\eta$ is the symmetric mass ratio, $a$ is the Keplerian semimajor axis, and $e$ is the Keplerian eccentricity. We then introduce the 1PN-corrected orbital quantities to supplement those from Sec. A.1:
$P$ $=2\pi\sqrt{\frac{a^{3}}{M}}\left[1+\frac{(9-\eta)}{2}\frac{M}{a}\right]\,,$ (25)
$t$ $=\frac{P}{2\pi}[u(t)-e_{t}\cos u(t)-\ell]$ (26)
@@
$=\frac{2\pi}{P(1-e_{t}\cos u(t))}\frac{\tilde{h}}{\tilde{a}^{2}}\frac{-2\tilde{e}\sin u(t)(1-\tilde{e}\cos u(t))}{(1-\tilde{e}\cos u(t))^{4}}\,.$ (28)
To get from the first to the second line of Eq. (28), we compute $du/dt$ by differentiating Eq. (26). Between the fourth and fifth lines, we multiply by $1=(1-\tilde{e}\cos u(t))/(1-\tilde{e}\cos u(t))$ to get an even power of $(1-\tilde{e}\cos u(t))$ in the denominator. To this 1PN order, we apply $(1-\tilde{e}\cos u(t))^{2}=(1-e_{t}\cos u(t))(1-e_{\phi}\cos u(t))$. Thus, Eq. (28) simplifies to
@@ A.3 Deriving the mapping from $\theta_{\rm orb}$ to $\theta_{\rm ID}$
Again, we follow Ref. Poisson and Will (2014) for deriving a mapping between the orbital parameters $\vec{\theta}_{\rm orb}=(a,e,\ell)$ and the initial data parameters $\vec{\theta}_{\rm ID}=(\Omega,v_{r},D)$.
@@
The $a$ and $e$ in Eq. (32) are the Newtonian parameters for semimajor axis and eccentricity, respectively.
@@
Finally, we consider the angular velocity, $\Omega(t)$. We begin with the first line of Eq. (28) for the angular velocity
@@
Then, we substitute Eqs. (20) and (27) to achieve
@@
Further substituting the 1PN approximations of Eqs. (27), (20), (21) yields the final expression
@@ A.4 Spin update derivations
First, we define $R_{\rm BH\to in}(t=t_{T})$, the rotation from the corotating BH frame into the inertial frame. The target spins $\vec{\chi}_{A,T}(t=t_{T})$ are defined in the BH frame which might not coincide with the inertial frame at $t_{T}$. The rotation matrix $R_{\rm BH\to in}(t=t_{T})$ which transforms from the BH to the inertial frame at $t_{T}$, is formed through the basis vectors $\hat{n}(t_{T}),\hat{\lambda}(t_{T})$, and $\hat{L}(t_{T})$, defined in the inertial frame. Again, we utilize the inertial frame vector notation from Ref. Buonanno et al. (2011):